OpenStudy (anonymous):

let f(x)= 2x-6 and g(x)= -3x-7. find (FoG)(3)

OpenStudy (anonymous):

(FoG)(3)=F(G(3))=G(3)=-3(3)-7=-16 So F(-16)=2(-16)-6=-38